The orbit of the earth around the sun is an ellipse. The sun is at one of the foci of this ellipse. If the length of the major axis of ellipse is 300 million km and the eccentricity is 0.0167, find the minimum and maximum distance of the earth from the sun.

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
Max. distance is when earth is at the far end of major axis = a + c = a(1+e) = 300 (1.0167) million kms
Min. distance is when earth is at the near end of major axis= a - c = a(1-e) 300(1-0.0167) million kms

Thanks
bharat
askiitians faculty
IIT Delhi